OpenStudy (anonymous):

El capitan in Yosemite national park is 3593 ft high. How long would it take a carabiner to fall freely from the top?

OpenStudy (anonymous):

Does he accelerate as he falls?

OpenStudy (anonymous):

IF don't consider the drag it is t\[\sqrt{\frac{2h}{g}}\]

OpenStudy (anonymous):

It doesnt say

OpenStudy (anonymous):

It does.

OpenStudy (anonymous):

Any idea how fast

OpenStudy (anonymous):

Assuming the accleration to be 9.8ms\s^2 we get t =14.9489 seconds.But,first you need to convert ft into meters.

OpenStudy (anonymous):

Aron just told you that.

OpenStudy (anonymous):

3593 ft = 1095 m

OpenStudy (anonymous):

Aron g=32 feet per second

OpenStudy (anonymous):

You get the same answer!

OpenStudy (anonymous):

Aron i solved you equilateral triangle question it is 3/8 right